OpenStudy (simoner):

Find the Taylor polynomial T3(x) for the function f centered at the number a.

OpenStudy (simoner):

f(x)=1/x a=4

OpenStudy (amistre64):

whats the process?

OpenStudy (simoner):

i thought the answer was 1/4-(x-4)/16+2/64(x-4)^2-2/64(x-4)^3 but that was wrong

OpenStudy (simoner):

nevermind i figured it out

OpenStudy (amistre64):

f = x^(-1) f' = -x^(-2) f'' = 2x^(-3) 4^(-1) -4^(-2)(x-4) +2(4)^(-3)(x-4)^2 T3 is 3 terms or degree 3?